How many grams are there in 0.0457 moles of (NH4)2S?
1
Expert's answer
2018-04-04T04:25:54-0400
n((NH4)2S) = 0.0457 moles

M((NH4)2S) = 14*2 + 1*8 + 32 = 68 g/mol

m((NH4)2S) = n*M = 0.0457mol*68g/mol = 3.1076 g